PartnerObserving Log for 2007 April 16/17

Observers: J. Harris, D. Sand
Instrument Support:
Telescope Operator:
Instrument: LBC-Blue

$ 19:00: Camera shutter being examined

$ 23:15: They didn't find anything wrong with the shutter; it passes all tests and is moving freely. They're going to put it all back together, and we may be able to observe shortly (although the humidity is currently 87%...)

$ 00:20: Attempting to bring camera online.

$ 00:30: Taking some test dome flats. The problem is, that if the shutter is stuck open, then flats won't reveal it. We need to compare darks, biases and dome flats. A 12-second "dark" with the B filter looks like a dark, while a 12-sec exposure with the B filter looks like a dome flat. So, the shutter must be working.

We're waiting for the humidity to drop below 90%, then we can open. Also having some trouble bringing the mirror up. We also can't get the dome shutter to open.

$ 02:30: They found that a gear on the dome shutter was broken, it will need to be fixed in the daytime.

Good night Safford!

-- JasonHarris - 17 Apr 2007
Topic revision: r5 - 21 Jul 2009, NormCushing
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback